A student and a professor each choose a number between 1 and 6 (1 and 6 are both possible choices). What is the probability that the two choose the same number?
Answer by stanbon(75887)

A student and a professor each choose a number between 1 and 6 (1 and 6 are both possible choices). What is the probability that the two choose the same number?
----------------------------
There are 6*6 = 36 possible pairs
---
There are 6 pairs that have the same number: 1/1,2/2,3/3 etc.
----
P(pair of same digits) = 6/36 = 1/6
----
Another way to think of the problem.
The student chooses any of the 6 digits.
Prob the professor chooses the same digit = 1/6
=================================================
Cheers,
Stan H.